40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 20 12:55:02 2019, gautam, Update, CDS, c1iscaux testing 
    Reply  Mon Sep 23 10:49:34 2019, rana, Update, CDS, c1iscaux testing 
    Reply  Wed Sep 25 20:10:13 2019, Koji, Update, CDS, c1iscaux testing testDetectMons_190925.pngtestIPPOS_190925.png
       Reply  Thu Sep 26 20:09:40 2019, Koji, Update, CDS, c1iscaux testing Wht1.pdfWht2.pdfWht3.pdfWht4.pdflock.png
          Reply  Fri Sep 27 15:59:53 2019, gautam, Update, CDS, c1iscaux testing 
          Reply  Wed Oct 2 01:11:40 2019, Koji, Update, CDS, c1iscaux testing 8x
             Reply  Fri Oct 4 01:57:09 2019, Koji, Update, CDS, c1iscaux testing P_20191003_172956_vHDR_On.jpgTF.pdfPSD.pdf191003_AA_Filter.zip
                Reply  Sat Oct 5 00:03:21 2019, Koji, Update, CDS, c1iscaux testing REFL1_GAIN1.pdfREFL1_GAIN2.pdfREFL2_GAIN1.pdfREFL2_GAIN2.pdf
                   Reply  Tue Oct 8 03:32:42 2019, Koji, Update, CDS, CM servo board testing 7x
                      Reply  Tue Oct 8 17:59:29 2019, Koji, Update, CDS, CM servo board testing (portal) 6x
                         Reply  Tue Oct 8 18:42:39 2019, Koji, Update, CDS, CM servo board testing boosts.pdf
                            Reply  Mon Oct 14 16:06:28 2019, Koji, Update, CDS, CM servo board testing pole_zero_filter.pdf
                               Reply  Mon Oct 14 16:19:30 2019, Koji, Update, CDS, CM servo board testing testb2.pdf
                                  Reply  Mon Oct 14 16:25:03 2019, Koji, Update, CDS, CM servo board testing servo_out.pdf
                                     Reply  Mon Oct 14 16:34:42 2019, Koji, Update, CDS, CM servo board testing in12_output_offset.pdfin12_input_offset.pdfin12_input_offset2.pdf
                                        Reply  Mon Oct 14 17:32:28 2019, Koji, Update, CDS, Portal Elog entry for the recent CM servo board tests CM_Servo_Diagram.png
                         Reply  Tue Oct 8 20:23:03 2019, gautam, Update, CDS, c1iscaux testing 
    Reply  Mon Sep 30 11:20:43 2019, gautam, Update, CDS, c1iscaux testing - CM board code updated CMsoftTest.png
       Reply  Mon Sep 30 15:51:59 2019, gautam, Update, CDS, c1iscaux - some admin 
Message ID: 14903     Entry time: Fri Sep 20 12:55:02 2019     Reply to this: 14905   14906   14912
Author: gautam 
Type: Update 
Category: CDS 
Subject: c1iscaux testing 

I was hoping that the dark / electronics noise level on the LSC photodiodes would be sufficient for me to test the whitening gain switching on the iLIGO Pentek whitening boards. However, this does not seem to be the case. I guess to be thorough, we have to do this kind of test. It's a bit annoying to have to undo and redo the SMA connections, but I can't think of any obvious easier way to test this functionality. More annoyingly, the sensing matrix infrastructure necessary to do the kind of test described in the linked elog is only available for some PDs. I don't really want to modify the c1cal model and go through another mass reboot cycle.

While I was at it, I was also thinking about the tests we want to do. Here is a quick first pass - if you can think of other tests we ought to do, please add them to the list!

  1. Whitening gain switching on the D990694 boards.
    • Need to inject some signal to do this in a clean way. 
    • With some signal injected, we need to switch the whitening gain through the 15 available levels and confirm that we see a +3dB gain for each step.
    • An example script to do this operation and make a diagnostic plot is at /cvs/cds/caltech/target/c1iscaux/testScripts/testWhtGain.py.
  2. AA enable/disable on the D000076 boards. Do we really need this functionality? Can't we permanently enable the AA, as was done for WF2?
    • Need to measure the TF with an SR785 or drive a high-freq line and confirm that the aliased peak height is attenuated as expected in DTT.
  3. LO Det Mon channel check
    • Zeroth level test can be done by turning Marconi OFF/ON, and confirming we see a change in the corresponding monitor channel, like I did here.
    • A more rigorous diagnostic would require these channels to be calibrated to dBm of LO power.
  4. PD INTF board check
    • Zeroth level check can be done by shining light onto PDs one at a time and confirming that the correct channel shows a response.
    • A more rigorous diagnostic would require these channels to be calibrated to mW of optical power incident on the PDs.
  5. QPD INTF board check
    • This is the IP-POS QPD readback.
    • Need to confirm the quadrant mapping, and that Pitch is really Pitch, Yaw is really Yaw.
    • A more rigorous diagnostic would require these channels to be calibrated to mm of position shift.
  6. CM Board
    • Need to determine what tests need to be done.
    • I have not yet implemented the fix for the MBBO gain channels for all the gains - only REFL1_GAIN is set up correctly now. Need to look at the hardware for the correct addressing of bits.
  7. ALS INTF board
    • This board isn't actually connected yet, pending strain relief of cabling at 1Y2.
    • The calibration of the board output volts to dBm is known, so we can easily check this functionality.
ELOG V3.1.3-